# Unseen from Seen: Rewriting Observation-Instruction Using Foundation Models for Augmenting Vision-Language Navigation

### v1

@@ IV-C Ablation Study
We conduct ablation studies to analyze the effects of different method components and the impacts of adding our augmentation data during different training phases. We use CLIP ViT B/16 features for the ablation experiments. For the baseline in Table II, Fig. 4, and Table VII, we reproduce the baseline model DUET [39] with CLIP VIT B/16 visual features. In Table VI, since we aim to verify the effectiveness of our rewritten data and exclude the effect of high-quality human annotated instructions, we train the baseline using the widely used R2R augmentation dataset [74], which can provide a fair comparison of different kinds of augmentation data.
Figure 5: Visualization examples of rewritten object-enriched scene descriptions, generated panorama, extracted ground-truth sequence from the generated panorama, and the rewritten instruction. Generated new scene objects and modality-aligned objects in instructions are denoted in red and blue fonts/boxes, respectively.
TABLE VIII: Comparison with the baseline method [39] under the low-resource settings. Using CLIP ViT B/16 as image features.
@@
Figure 6: Visualization comparison of instructions generated by the Speaker model [53] and our RAM. Modality-aligned objects in the instruction are denoted in blue fonts/boxes.
@@ IV-E Visualization
Rewritten Observation-Instruction Examples. Fig. 5 presents a visualization example of rewritten data by our RAM. From the rewritten object-enriched descriptions generated by the LLM, we can find that diverse objects, e.g., armchair and coffee table, are indicated explicitly. Benefiting from it, the panorama generated by the T2IM contains these newly introduced objects with new spatial layouts, leading to great diversity enhancement of the environment. Moreover, we can observe that the extracted single-view images are also high-quality compared to the images taken by the camera in Matterport3D dataset. We can also find that the rewritten instruction successfully indicates modality-aligned objects, e.g., hallway and plants in the new trajectory, revealing the effectiveness of our observation-contrast instruction rewriting.
@@ II-A Scale of augmented data
For each dataset R2R, REVERIE, and R4R, the augmentation data are 3 times larger than the original dataset, shown in Table IX, which is significantly smaller than other methods such as AutoVLN [9] and ScaleVLN [10]. All our augmentation data are utilized in both pretraining and finetuning phases.
@@ II-B Choices of foundation models and implementation details
For generating observations, the reason for selecting Multidiffusion [60] as the T2I model used in our pipeline has been indicated in our original manuscript, i.e., the Multidiffusion model is more suitable for generating panoramas. In the original Multidiffusion [60] paper, the authors have also demonstrated its superior capability of panorama generation over Stable Diffusion.
For generating scene descriptions, we have compared the descriptions obtained through different popular VLMs, including BLIP [75], Tag2text [68], and LLaVA [76], and chose Tag2Text [68] finally since it tends to capture informative and salient objects in the observation image. We also put a comparison visualization example here in Fig. 8. The results in Fig. 8 reveal that BLIP tends to miss critical objects in the scene and LLaVA often generates redundant or irrelevant information for image generation. In contrast, Tag2text can accurately identify noteworthy objects and generate informative captions.
@@
1 Observation Descriptions VLM - $<$ 1 hours
2 Rewritten Observation Descriptions LLM 10 dollars $<$ 30 minutes
@@ 3 Augmented Observations T2I Model - $\sim$ 30 hours
4 Rewritten Instructions LLM 10 dollars $<$ 30 minutes
Figure 8: Captions of single-view images and Scene descriptions of panoramas obtained by different VLMs [75, 68]. Scene objects are denoted in red fonts/boxes, and the wrong noun in the caption is denoted in blue box.
@@ 4 Rewritten Instructions LLM 10 dollars $<$ 30 minutes
The VLM, LLM, and T2IM used in our data augmentation pipeline are all employed in an off-the-shelf manner. For the VLM (i.e., Tag2text [68]), we keep the original parameter setting (e.g. threshold for tagging and input image size) for obtaining scene descriptions. For LLM (i.e., ChatGPT [44]), we set the parameter temperature to 0.8 and presence_penalty to 0. For the T2IM (i.e., Multidiffusion [60]), we fix the parameter num_inference_steps of the MultiDiffusion model to 30 for generating panoramas with high quality and resolution in a tolerable time consumption.
@@ II-C Details on the time, computational cost and resources
Details on the time, computational cost, and resources required for generating the augmented data by our RAM is presented in Table X. Concretely, we utilized VLM and T2I models to obtain observation descriptions and augmented observations on 8 NVIDIA RTX 3090 GPUs by splitting data into plenty of batches in 1 hour and 30 hours, respectively. Rewritten observation descriptions and rewritten instructions were collected by querying LLMs within 30 minutes. Overall, the entire data augmentation pipeline requires less than $\sim$2 days and relatively low computational resources.
TABLE XI: Comparison results with the baseline HAMT [38] on R2R dataset. †: using CLIP ViT B/16 as image features.
@@ III-A Performance comparison with baseline HAMT on R2R
We have incorporated a new baseline, HAMT [38], along with corresponding experiments on the R2R dataset, as presented in Table XI here. From Table XI, we can observe that our RAM outperforms the baseline HAMT in both Val Seen and Val Unseen, e,g, it brings $\sim$3improvements in SPL in Val Unseen. The results on different baselines sufficiently show the effectiveness and generalization ability of our proposed method.
Figure 9: Rewritten descriptions with different prompts and panoramas generated by T2I model with the corresponding rewritten descriptions. On the right side, we utilize the prompt adding object attribute constraints to generate rewritten description. Added objects in rewritten scene descriptions and generated panoramas are denoted in red fonts/boxes.
@@ III-B Comparison with ScaleVLN on the same data scale
Table XII shows a new ablation experiment with CLIP VIT B/16 visual features. Considering that acquiring data at the scale of ScaleVLN [10] is highly labor-intensive and poses significant storage challenges, we conduct an alternative experimental comparison between ScaleVLN [10] and our RAM. Specifically, we train DUET [39] agent using the ScaleVLN data with the same amount of augmented data as our RAM, i.e., we randomly sample 14,025 trajectory-instruction pairs (x3 times the data amount of the original R2R data) from the ScaleVLN [10] dataset in R2R dataset  [^2] The CLIP VIT B/16 visual features are provided by the ScaleVLN GitHub repository, https://github.com/wz0919/ScaleVLN..We have presented the result here in Table XII. From Table XII, we can observe that the “DUET+ScaleVLN(subset)” setting surpasses the baseline, yet its performance is inferior to that of our RAM. Moreover, this setting’s performance on the Val Seen drops by approximately $\sim$6in SR and $\sim$8in SPL compared to the baseline, which may be attributed to the introduction of a large amount of unfamiliar real-world simulator data causing the agent to lose its generalization capacity in the original simulation environment. Additionally, we can see that our RAM outperforms ScaleVLN in Val Unseen with the same data scale, which further validates the high quality of our augmented data and the effectiveness of our proposed mixed training mechanism.
TABLE XII: Comparison with baseline method [39] and ScaleVLN [10] with the same scale as ours (x3 times larger than the R2R dataset). †: using CLIP ViT B/16 as visual features.
@@ III-C Comparison with object deletion
TABLE XIII: Ablation results on object addition and object deletion. We conduct this experiment using CLIP ViT B/16 as visual features.
@@ III-D Controlled object generation
To investigate the potential of our RAM in controlled object generation for creating specific VLN scenarios, we incorporate the object attribute constraints on the LLM prompt presented in Fig. 9. From Fig. 9, we can observe that by adding the constraints, the LLM can successfully generate objects with desired styles, e.g., the “bar stools” and “plush sofa” which have a moderate size. And the T2IMs also successfully generate the desired objects in the panoramas. These results reveal the potential of extending our method to generate VLN scenarios with specific styles.
@@
[…]
@@
11:        Use an LLM to extract the sequential landmarks $U=\{U_{k}\}_{k=1}^{M}$ from the original instruction $I$

### v2

@@ IV-C Ablation Study
We conduct ablation studies to analyze the effects of different method components and the impacts of adding our augmentation data during different training phases. We use CLIP ViT-B/16 features for the ablation experiments. For the baseline in Table II, Fig. 4, and Table VII, we reproduce the baseline model DUET [39] with CLIP VIT-B/16 visual features. In Table VI, since we aim to verify the effectiveness of our rewritten data and exclude the effect of high-quality human annotated instructions, we train the baseline using the widely used R2R augmentation dataset PREVALENT [78], which can provide a fair comparison of different kinds of augmentation data.
Figure 5: Visualization examples of rewritten object-enriched scene description, generated panorama, extracted ground-truth sequence from the generated panorama, and the rewritten instruction. Newly generated scene objects and modality-aligned objects in the instruction are denoted in red boxes and bold fonts, respectively.
TABLE VIII: Comparison with the baseline method [39] under the low-resource settings. Using CLIP ViT-B/16 as image features.
@@
Figure 6: Visualization comparison of instructions generated by the Speaker model [58] and our RAM. Modality-aligned objects in the instructions are denoted in red boxes and bold fonts.
@@ IV-E Visualization
Rewritten Observation-Instruction Examples. Fig. 5 presents a visualization example of rewritten data by our RAM. From the rewritten object-enriched descriptions generated by the LLMs, we can find that diverse objects, e.g., armchair and coffee table, are indicated explicitly. Benefiting from it, the panorama generated by the T2IM contains these newly introduced objects with new spatial layouts, leading to great diversity enhancement of the environment. Moreover, we can observe that the extracted single-view images are also high-quality compared to the images taken by the camera in Matterport3D dataset. We can also find that the rewritten instruction successfully indicates modality-aligned objects, e.g., hallway and plants in the new trajectory, revealing the effectiveness of our observation-contrast instruction rewriting.
@@ II-A Scale of augmented data
For each dataset (R2R, REVERIE, and R4R), the augmented data are three times larger than the original dataset, shown in Table IX, which is significantly smaller than other methods such as AutoVLN [9] and ScaleVLN [10]. All our augmentation data are utilized in both pretraining and finetuning phases.
@@ II-B Choices of foundation models and implementation details
For generating observations, the reason for selecting Multidiffusion [65] as the text-to-image model used in our pipeline has been indicated in our original manuscript, i.e., the Multidiffusion model is more suitable for generating panoramas. In the original Multidiffusion [65] paper, the authors also demonstrated its superior capability of panorama generation compared to Stable Diffusion.
For generating scene descriptions, we have compared the descriptions obtained through different popular VLMs, including BLIP [79], Tag2Text [72], and LLaVA [80], and finally chose Tag2Text [72] since it tends to capture informative and salient objects in the observation image. We also put a comparison visualization example here in Fig. 8. The results in Fig. 8 reveal that BLIP tends to miss critical objects in the scene and LLaVA often generates redundant or irrelevant information for image generation. In contrast, Tag2Text can accurately identify noteworthy objects and generate informative captions.
@@
1 Observation Descriptions VLM - << 1 hours
2 Rewritten Observation Descriptions LLM 10 dollars << 30 minutes
@@ 3 Augmented Observations T2I Model - $\sim$ 30 hours
4 Rewritten Instructions LLM 10 dollars << 30 minutes
Figure 8: Captions of single-view images and scene descriptions of panoramas obtained by different VLMs [79, 72]. Scene objects are denoted in red fonts/boxes, and the incorrect noun in the caption is denoted in blue box.
@@ 4 Rewritten Instructions LLM 10 dollars << 30 minutes
The VLM, LLM, and T2IM used in our data augmentation pipeline are all employed in an off-the-shelf manner. For the VLM (i.e., Tag2Text [72]), we keep the original parameter setting (e.g. threshold for tagging and input image size) for obtaining scene descriptions. For the LLM (i.e., ChatGPT [49]), we set the parameter temperature to 0.8 and presence_penalty to 0. For the T2IM (i.e., MultiDiffusion [65]), we fix the parameter num_inference_steps of the MultiDiffusion model to 30 for generating panoramas with high quality and resolution in a tolerable time consumption.
@@ II-C Details on the time, computational cost and resources
Details on the time, computational cost, and resources required for generating the augmented data by our RAM are presented in Table X. Concretely, we utilized VLM and T2I models to obtain observation descriptions and augmented observations on 8 NVIDIA RTX 3090 GPUs by splitting data into multiple batches in 1 hour and 30 hours, respectively. Rewritten observation descriptions and rewritten instructions were collected by querying LLMs within 30 minutes. Overall, the entire data augmentation pipeline requires less than $\sim$2 days and relatively low computational resources.
TABLE XI: Comparison results with the baseline HAMT [38] on R2R dataset. †: using CLIP ViT-B/16 as image features.
@@ III-A Performance comparison with baseline HAMT on R2R
We have incorporated a new baseline, HAMT [38], along with corresponding experiments on the R2R dataset, as presented in Table XI here. From Table XI, we can observe that our RAM outperforms the baseline HAMT in both Val Seen and Val Unseen, e.g., it brings $\sim$3% improvements in SPL in Val Unseen. The results on different baselines demonstrate the effectiveness and generalization ability of our proposed method.
Figure 9: Rewritten descriptions with different prompts and panoramas generated by T2I model with the corresponding rewritten descriptions. On the right side, we utilize the prompt adding object attribute constraints to generate rewritten descriptions. Added objects in rewritten scene descriptions and generated panoramas are denoted in red boxes and bold fonts.
@@ III-B Comparison with ScaleVLN on the same data scale
[…]
@@ III-C Comparison with object deletion
TABLE XIII: Ablation results on object addition and object deletion. We conduct this experiment using CLIP ViT-B/16 as visual features.
@@ III-D Controlled object generation
To investigate the potential of our RAM in controlled object generation for creating specific VLN scenarios, we incorporate the object attribute constraints on the LLM prompt presented in Fig. 9. From Fig. 9, we can observe that by adding the constraints, the LLM can successfully generate objects with desired styles, e.g., the “bar stools” and “plush sofa” which have a moderate size. And the T2IM also successfully generates the desired objects in the panoramas. These results reveal the potential of extending our method to generate VLN scenarios with specific styles.
@@
7:   Use $\mathrm{Equirec2Perspec}$ algorithm to get single-view images, i.e., $\{O^{r}_{t,k}\}_{k=1}^{K}=\mathrm{Equirec2Perspec}\left(J_{inv},R,O^{r}_{t}\right)$
@@
11:   Use an LLM to extract the sequential landmarks $U=\{U_{k}\}_{k=1}^{M}$ from the original instruction II